# A quantitative Borg-Levinson theorem for a large class of unbounded potentials

### v2

@@ 1. Introduction
$\tilde{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Omega)}\right]$
@@ 2.2. Dirichlet to Neumann map
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. It follows from Theorem B.1 in Appendix B that, for all $\varphi\in W^{2-1/q,q}(\Gamma)$, the BVP (2.7) admits a unique solution $u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$. Furthermore, we have
(2.8) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\varkappa\|\varphi\|_{W^{2-1/q,q}(\Gamma)},$
@@
$\Lambda_{V}(\lambda):\varphi\in W^{2-1/q,q}(\Gamma)\mapsto\gamma_{1}u_{V}(\lambda)(\varphi)\in W^{1-1/q,q}(\Gamma).$
By the continuity of the trace map $w\in W^{2,q}(\Omega)\mapsto\gamma_{1}w\in W^{1-1/q,q}(\Gamma)$, it follows from (2.8) that the mapping $\Lambda_{V}(\lambda)$ defines a bounded operator.
@@ 2.3. Analyticity of solutions
By using that $\Delta\phi_{V}^{k}=(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{q}(\Omega)$, we get $\phi_{k}\in W^{2,q}(\Omega)$. In light of $W^{2,q}$ a priori estimate (e.g. [13, Theorem 2.3.3.6]), (2.10) yields
@@
Let $\lambda\in\rho(A_{V})$, $\varphi\in W^{2-1/q,q}(\Gamma)$ and $u:=u_{V}(\lambda)(\varphi)$. Then
$0=\int_{\Omega}(-\Delta u+(V-\lambda)u)\overline{\phi_{V}^{k}}dx.$
@@
(2.14) $u_{V}^{(j)}(\lambda)(\varphi)=-j!\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{(\lambda_{V}^{k}-\lambda)^{j+1}}\phi_{V}^{k},\quad j\geq 0.$
The following notation will be used hereinafter.
$w^{(j)}:=\frac{d^{j}}{d\lambda^{j}}w.$
Let $\mathscr{S}$ be the bounded affine function $f\in L^{q}(\Omega)\mapsto\mathscr{S}f:=w\in W^{2,q}(\Omega)$, where $w$ is the unique solution of the BVP
$\Delta w=f\;\mathrm{in}\;\Omega,\quad u=\varphi\;\mathrm{on}\;\Gamma.$
By observing that $u_{V}(\lambda)(\varphi)=\mathscr{S}((\lambda-V)u_{V}(\lambda)(\varphi))$, we conclude that $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$ is holomorphic and by induction in $j\geq 1$, we have, where $u(\lambda):=u_{V}(\lambda)(\varphi)$,
(2.15) $(-\Delta+V-\lambda)u^{(j)}(\lambda)=ju^{(j-1)}(\lambda)\;\mathrm{in}\;\Omega,\quad u^{(j)}(\lambda)=0\;\mathrm{on}\;\Gamma.$
@@ 2.5. The difference between two Dirichlet to Neumann maps
$\mathcal{B}=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$
@@
$\|\varphi\|:=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
@@
Recall also that $\mathcal{B}:=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$ is equipped with the norm
$\|\varphi\|=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
@@
$\hskip 85.35826pt+\tau^{\sigma}\int_{0}^{1}\|\Lambda_{V_{1}}(\mu_{\tau}^{s}(\varphi))-\Lambda_{V_{2}}(\mu_{\tau}^{s}(\varphi))\|_{L^{q}(\Gamma)}ds.$
@@
(3.13) $\|\gamma_{1}v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|,$
@@
(3.14) $\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\tau^{n+2}\mathscr{D}(V_{1},V_{2})\|\varphi\|$
@@ Appendix B BVPs with unbounded lower order term
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. For all $f\in L^{q}(\Omega)$ and $\varphi\in W^{2-1/q,q}(\Gamma)$ the BVP (B.1) has a unique solution $u=u_{V}(\lambda)(\varphi,f)\in W^{2,q}(\Omega)$. In addition, there exists a constant $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ such that
(B.2) $\|u_{V}(\lambda)(\varphi,f)\|_{W^{2,q}(\Omega)}\leq\varkappa\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
@@
In this proof, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$ and $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ are generic constants.
Let $F\in W^{2,q}(\Omega)$ so that $\gamma_{0}F=\varphi$ and
$\|F\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{0}\|\varphi\|_{W^{2-1/q,q}(\Gamma)}.$
If $G=-(-\Delta+V-\lambda)F+f$ then $G\in L^{q}(\Omega)$ and
(B.3) $\|G\|_{L^{q}(\Omega)}\leq\mathbf{c}\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
By density, we can find a sequence $(G_{k})$ in $L^{2}(\Omega)$ converging in $L^{q}(\Omega)$ to $G$. Set for any $k$
(B.4) $v_{k}=R_{V}(\lambda)G_{k}.$
Then $v_{k}\in H_{0}^{1}(\Omega)$ and in consequence
$\Delta v_{k}=-Vv_{k}+\lambda v_{k}+G_{k}\in L^{q}(\Omega).$
By [13, Theorem 2.4.2.6, page 241], $v_{k}\in W^{2,q}(\Omega)$. Then we proceed as in the proof of Lemma 2.4 to obtain
$\|v_{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|(-\Delta+\lambda_{0})v_{k}\|_{L^{q}(\Omega)},$
@@
$\|v_{k}\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}\|-Vv_{k}+(\lambda+\lambda_{0})v_{k}+G_{k}\|_{L^{q}(\Omega)}$
$\leq\mathbf{c}\left(\|V\|_{L^{n/2}(\Omega)}\|v_{k}\|_{L^{2}(\Omega)}+|\lambda+\lambda_{0}|\|v_{k}\|_{L^{q}(\Omega)}+\|G_{k}\|_{L^{q}(\Omega)}\right).$
From (B.4), $(v_{k})$ is bounded in $H_{0}^{1}(\Omega)$. Hence the last inequalities show that $(v_{k})$ is bounded in $W^{2,q}(\Omega)$ with
(B.5) $\|v_{k}\|_{W^{2,q}(\Omega)}\leq\varkappa\|G_{k}\|_{L^{q}(\Omega)}.$
Here and henceforth, $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ is a generic constant.
Now as $W^{2,q}(\Omega)$ is reflexive, subtracting if necessary a subsequence, we may assume that $(v_{k})$ converges weakly in $W^{2,p}(\Omega)\cap W_{0}^{1,q}(\Omega)$ to $v\in W^{2,q}(\Omega)\cap W_{0}^{1,q}(\Omega)$. Then $\Delta v+Vv-\lambda v=G$ in the distributional sense and, by using that a norm is weakly lower semi-continuous, we obtain from (B.5)
(B.6) $\|v\|_{W^{2,q}(\Omega)}\leq\varkappa\|G\|_{L^{q}(\Omega)}.$
The function $u=F+v\in W^{2,p}(\Omega)$ is clearly a solution of the BVP (B.1) and inequality (B.2) is a straightforward consequence of inequalities (B.3) and (B.6).
The uniqueness of solutions of the BVP (B.1) follows from the fact that $\lambda$ is not an eigenvalue of $A_{V}$. ∎

### v3

@@ 1. Introduction
$\tilde{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,\tilde{q}}(\Omega)}\right],$
@@ 2.2. Dirichlet to Neumann map
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. It follows from Theorem B.1 in Appendix B that, for all $\varphi\in W^{2-1/p,p}(\Gamma)$, the BVP (2.7) admits a unique solution $u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$. Furthermore, we have
(2.8) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,p}(\Omega)}\leq\varkappa\|\varphi\|_{W^{2-1/p,p}(\Gamma)},$
@@
$\Lambda_{V}(\lambda):\varphi\in W^{2-1/p,p}(\Gamma)\mapsto\gamma_{1}u_{V}(\lambda)(\varphi)\in W^{1-1/p,p}(\Gamma).$
By the continuity of the trace map $w\in W^{2,p}(\Omega)\mapsto\gamma_{1}w\in W^{1-1/p,p}(\Gamma)$, it follows from (2.8) that the mapping $\Lambda_{V}(\lambda)$ defines a bounded operator.
@@ 2.3. Analyticity of solutions
By using that $\Delta\phi_{V}^{k}=(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{p}(\Omega)\subset L^{q}(\Omega)$ (with continuous embedding), we get $\phi_{k}\in W^{2,q}(\Omega)$. In light of $W^{2,q}$ a priori estimate (e.g. [13, Theorem 2.3.3.6]), (2.10) yields
@@
Let $\lambda\in\rho(A_{V})$, $\varphi\in W^{2-1/p,p}(\Gamma)$ and $u:=u_{V}(\lambda)(\varphi)$. Then
$0=\int_{\Omega}\left(\nabla u\cdot\nabla\overline{\phi_{V}^{k}}+(V-\lambda)u\overline{\phi_{V}^{k}}\right)dx,$
where we used that $\nabla u\in L^{2}(\Omega,\mathbb{C}^{n})$ and $u\in L^{p^{\prime}}(\Omega)$.
@@
(2.14) $u_{V}^{(j)}(\lambda)(\varphi)=-j!\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{(\lambda_{V}^{k}-\lambda)^{j+1}}\phi_{V}^{k},\quad j\geq 0,$
where we used the notation
$w^{(j)}:=\frac{d^{j}}{d\lambda^{j}}w,$
which we will keep from now on.
In fact, we have $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$ is holomorphic as shown by the following proposition.
Proposition 2.2.
For all $V\in\mathscr{V}$, $\varphi\in W^{2-1/p,p}(\Gamma)$, $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$ is holomorphic and $u^{(1)}(\lambda):=u_{V}^{(1)}(\lambda)(\varphi)$ is the solution of the BVP
(2.15) $(-\Delta+V-\lambda)u^{(1)}(\lambda)=u_{V}(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad u=0\;\mathrm{on}\;\Gamma,$
Proof.
Let $V\in\mathscr{V}$, $\lambda\in\rho(A_{V})$ and $\mu\in B(0,\mathrm{dist}(\lambda,\sigma(A_{V})/2))$. From Theorem B.1 and its proof, the constant $\mathbf{c}_{\lambda+\mu}$ in (B.5) can be chosen such that
$\mathbf{c}_{\lambda}:=\sup_{\mu\in B(0,\mathrm{dist}(\lambda,\sigma(A_{V})/2)}\mathbf{c}_{\lambda+\mu}<\infty.$
Since $v=u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)$ is the solution of the BVP
$(-\Delta+V-(\lambda+\mu))v=\mu u(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad v=0\;\mathrm{on}\;\Gamma,$
we obtain from (B.5)
(2.16) $\|u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}|\mu|\|u(\lambda)(\varphi)\|_{L^{p}(\Omega)}.$
Hence, $z\in\rho(A_{V})\mapsto u_{V}(z)(\varphi)\in W^{2,p}(\Omega)$ is continuous.
Next, let $u^{(1)}(\lambda)$ be the solution of (2.15) and set
$w:=u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)-\mu u^{(1)}(\lambda).$
We verify that $w$ is the solution of the BVP
$(-\Delta+V-\lambda)w=\mu[u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)]\;\mathrm{in}\;\Omega,\quad w=0\;\mathrm{on}\;\Gamma,$
Applying (B.5) and then (2.16), we obtain
$\|u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)-\mu u^{(1)}(\lambda)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}^{2}|\mu|^{2}\|u(\lambda)(\varphi)\|_{L^{p}(\Omega)}.$
This inequality leads to the expect result. ∎
Since $W^{2,p}(\Omega)$ is continuously embedded in $W^{2,p}(\Omega)$, the map $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$ is also holomorphic for all $\varphi\in W^{2-1/p,p}(\Gamma)$. This observation will be used in the following.
Let $V\in\mathscr{V}$, $\varphi\in W^{2-1/p,p}(\Gamma)$ and $\lambda\in\rho(A_{V})$. By induction in $j\geq 1$, we derive from (2.15)
(2.17) $(-\Delta+V-\lambda)u_{V}^{(j)}(\lambda)(\varphi)=ju_{V}^{(j-1)}(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad u^{(j)}(\lambda)=0\;\mathrm{on}\;\Gamma.$
@@ 2.5. The difference between two Dirichlet to Neumann maps
$\mathcal{B}=W^{2-1/p,p}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$
@@
$\|\varphi\|:=\|\varphi\|_{W^{2-1/p,p}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
@@
Recall also that $\mathcal{B}:=W^{2-1/p,p}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$ is equipped with the norm
$\|\varphi\|=\|\varphi\|_{W^{2-1/p,p}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
@@
$\hskip 85.35826pt+\tau^{\sigma}\int_{0}^{1}\|\Lambda_{V_{1}}^{(1)}(\mu_{\tau}^{s}(\varphi))-\Lambda_{V_{2}}^{(1)}(\mu_{\tau}^{s}(\varphi))\|_{L^{q}(\Gamma)}ds.$
@@
(3.13) $\|\gamma_{1}v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
@@
(3.14) $\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\tau^{n+2}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
@@ Appendix B BVPs with unbounded lower order term
We will use the following lemma, the proof of which is quite similar to that of Lemma 2.2.
Lemma B.1.
Let $V\in\mathscr{V}$, $\lambda\in\rho(A_{V})$. Then
(B.1) $\|R_{V}(\lambda)f\|_{L^{2}(\Omega)}\leq\frac{1}{\mathrm{dist}(\lambda,\sigma(A_{V}))}\|f\|_{L^{2}(\Omega)},\quad f\in L^{2}(\Omega),$
(B.2) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|f\|_{L^{2}(\Omega)},\quad f\in L^{2}(\Omega),$
(B.3) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|f\|_{L^{p}(\Omega)},\quad\quad f\in L^{p}(\Omega).$
Here, $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)>0$ is a constant.
@@
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. For all $f\in L^{p}(\Omega)$ and $\varphi\in W^{2-1/p,p}(\Gamma)$ the BVP (B.4) has a unique solution $u=u_{V}(\lambda)(\varphi,f)\in W^{2,p}(\Omega)$. In addition, there exists a constant $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)>0$ such that
(B.5) $\|u_{V}(\lambda)(\varphi,f)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}\left(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}\right).$
@@
In this proof, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$, $\mathbf{c}=\mathbf{c}(n,\Omega,V)>0$ and $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)$ are generic constants.
Let $F\in W^{2,p}(\Omega)$ so that $\gamma_{0}F=\varphi$ and
$\|F\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{0}\|\varphi\|_{W^{2-1/p,p}(\Gamma)}.$
If $G=-(-\Delta+V-\lambda)F+f$ then $G\in L^{p}(\Omega)$ and
(B.6) $\|G\|_{L^{p}(\Omega)}\leq\mathbf{c}_{\lambda}\left(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}\right).$
It follows from (B.3) that $v=R_{V}(\lambda)f\in H_{0}^{1}(\Omega)$ and
(B.7) $\|v\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|G\|_{L^{p}(\Omega)}.$
Furthermore, we have
$\Delta v=-Vv+\lambda v+G\in L^{p}(\Omega).$
By [13, Theorem 2.4.2.6, page 241], we obtain $v\in W^{2,p}(\Omega)$. Then we proceed as in the proof of Lemma 2.4 to obtain
$\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}\|(-\Delta+\lambda_{0})v\|_{L^{p}(\Omega)},$
@@
$\|v\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}_{0}\|-Vv+(\lambda+\lambda_{0})v+G\|_{L^{p}(\Omega)}$
$\leq\mathbf{c}_{0}\left(\|V_{0}\|_{L^{n/2}(\Omega)}\|v\|_{L^{p^{\prime}}(\Omega)}+|\lambda+\lambda_{0}|\|v\|_{L^{p}(\Omega)}+\|G\|_{L^{p}(\Omega)}\right).$
The last inequality, combined with (B.7), gives
$\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}\|G\|_{L^{p}(\Omega)}.$
That is, we have
(B.8) $\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}).$
We verify that $u=F+v\in W^{2,p}(\Omega)$ is a solution of the BVP (B.4) and inequality (B.5) follows from (B.8). This solution is unique because $\lambda$ is not an eigenvalue of $A_{V}$. ∎